Break-Even Sales and Sales Mix for a Service Company Zero Turbulence Airline provides air transportation services between Los Angeles, California, and Kona, Hawall. A single Los Angeles to Kona round-trip flight has the operating statistics: Fuel Flight crew salaries Airplane depreciation Variable cost per passenger-business dass Variable cost per passengermeconomy class Round-trip ticket price-business class Round-trip ticket price economy class $7.000 3,200 3,480 140 120 800 300 It is assumed that the fuel, crew salaries, and äirplane deprediation are fixed, regardless of the number of seats sold for the round-trip flight. is 10% business dass a compute the break even number o seats sold on a single round-trip flight for the overall product E. Assume that the overall product m economy dlass tickets. Total number of seats at break-even b. How many business class and economy class seats would be sold at the breakleus Business class seats at break-even Economy class seats at break-even seats seats seats
Answer #1

Answer

a ) calculate the contribution margin

unit selling price = 10% * 800 + 90% *300

= 350

unit variable price = 10% * 140 + 90% * 120

= 122

= 350 - 122 = 228

break even point = total fixed costs / contribution margin unit

= 13860 /228

= 60

b calculate the business class and economy class as follow

business class = 10% * 60 = 6

economy class = 90% *60 = 54

total = 60